OpenStudy (anonymous):

A custom t shirt company charges a $200 set up fee and $5 per t shirt. Approximately how may t shirts must be produced so that the total cost per t shirt is $5.33?

OpenStudy (anonymous):

answer options are a.1,000 tshirts b.200 tshirts c.600 tshirts d.1,266

OpenStudy (anonymous):

help me guys please :(

OpenStudy (displayerror):

Let \(x\) represent the number of shirts you sell. The cost per shirt would then be given by \[\frac{\left(200+5x\right) \ \text{dollars}}{x \ \text{shirts}}\] ^ Look at the units of the fraction above. Reading from top to bottom, you get dollars per shirt (this is to verify that the fraction we have gives us the value we want). How much should the cost per shirt equal? We're told that in the question statement: \($5.33\). This means we can set the fraction above equal to \($5.33\): \[\frac{200+5x }{x} = $5.33\] Now if you solve the fraction above, you should get \(x\), the number of shirts sold.

OpenStudy (anonymous):

i got 606

OpenStudy (displayerror):

Yep! Now which choice is 606 closest to? That should be your answer.

OpenStudy (anonymous):

thank you